# Cusp bifurcations: numerical detection via two-parameter continuation and computer-assisted proofs of existence

### v1

@@ 1 Introduction
A cusp bifurcation is a codimension-two bifurcation that provides an organizing center at which two saddle-node bifurcations come together. This fundamental phenomenon in bifurcation theory often gives rise to a hysteresis loop, which leads to the important notion of bistability. Besides its undeniable importance in dynamical systems theory, studying rigorously the occurrence of a cusp bifurcation is hindered by two main factors. First, as a codimension-two bifurcation, detecting a cusp requires varying two parameters. Second, actually proving the existence of a cusp in a system of nonlinear differential equations is nontrivial, as it requires having a precise location of the cusp point and then evaluating carefully high order derivatives at the point. In this paper, we propose an approach which addresses these issues. To numerically detect a cusp, we perform a two-parameter continuation of equilibria (e.g. see [4, 3, 21]) along which we apply the theory of Poincaré index to search for zeros of a specific two dimensional vector field (see Section 2). Once a cusp has been numerically detected during the continuation, we apply Newton’s method to a carefully chosen map (the cusp map defined in (14)) to refine the approximation which is then validated using a computer-assisted proof based on Newton-Kantorovich theorem (see Section 3). Finally, using the explicit control of the location of the cusp, we obtain rigorously the normal form on the center manifold. To formalize the ideas, let us introduce some notation and background.
@@
The process begins with the detection step (see Section 2), which is done by computing the Poincaré index of curves (i.e. the boundaries of two-dimensional simplices used in building a triangulation of a manifold $\mathcal{M}\subset\mathcal{E}$) with respect to a carefully chosen two-dimensional vector field. The computation of the Poincaré index requires in turn to perform the continuation of the singular value decomposition of $D_{x}f(x,\lambda)$ along the boundaries of simplices. If the Poincaré index is nontrivial, then a cusp bifurcation point has been detected “inside” the simplex. This approach is different from the more standard cusp detection, which consists of the two following steps: (a) do a one-dimensional continuation (e.g. pseudo-arclength continuation, see [27]) to detect a saddle-node bifurcation (fold); and (b) free an extra parameter, perform a continuation of folds, monitor the coefficient $b$ defined in (3), and claim that a cusp bifurcation has been detected when $b$ changes sign along the branch of folds. This well-known method can be applied using standard one-dimensional continuation softwares such as AUTO [20] and MATCONT [14]. In contrast, our approach relies on a two-dimensional continuation [4, 3, 21], and has the advantage of locating cusp bifurcation points on two-dimensional manifolds $\mathcal{M}\subset\mathcal{E}$. An obvious drawback is the computational overhead with respect to the approach based on one-dimensional continuations. However, a clear upside is that we construct explicitly a portion of the two-dimensional manifold $\mathcal{M}\subset\mathcal{E}$, as opposed to just one dimensional branches. In short, we get more at a larger cost.
@@ 2.2.1 Poincaré index
Let $g:(s_{1},s_{2})\in D\mapsto g(s_{1},s_{2})\in\mathbb{R}^{2}$ be continuous, with $D$ a bounded open subset of $\mathbb{R}^{2}$, and suppose $g$ has only isolated zeros in $D$. Let $\Gamma\subset D$ be a Jordan curve passing through no zeros of $g$. Let $\Delta\theta$ be the total change in the angle $\theta$ that $g(s_{1},s_{2})$ makes with respect to some fixed direction of $\mathbb{R}^{2}$ as $(s_{1},s_{2})$ traces $\Gamma$ once with positive (i.e. counterclockwise) orientation. Then, the Poincaré index of $\Gamma$ with respect to $g$ is defined as $\Delta\theta/2\pi$, and will be denoted by $\mathcal{I}_{g}(\Gamma)$.
Clearly, the number $\mathcal{I}_{g}(\Gamma)$ is an integer. Its importance lies in the following fundamental results.
@@
If $\Gamma\subset D$ is a Jordan curve containing no zeros of $g$ on it or in its interior, then $\mathcal{I}_{g}(\Gamma)=0$.
@@
Let $g:(s_{1},s_{2})\in D\mapsto g(s_{1},s_{2})\in\mathbb{R}^{2}$ be continuous, $D\subset\mathbb{R}^{2}$ open and bounded. Let $\Gamma\subset D$ be a Jordan curve passing through no zeros of $g$. If $\mathcal{I}_{g}(\Gamma)\neq 0$, then $g$ has a zero inside the region bounded by $\Gamma$.
@@ 2.2.2 Singular value decomposition of matrix valued functions
It is important to realize that the assumption of Theorem 2.3 is not at all restrictive but, rather, what one should generally expect to hold for a one parameter matrix valued function. In fact, one can show (again, see [15]) that the set of real ${n\times n}$ matrices having repeated singular values has dimension $n^{2}-2$, and therefore a generic [^3] We refer to [25] for necessary background on the concept of genericity $\mathcal{C}^{k}$ path $t\in\mathbb{R}\mapsto A(t)\in\mathbb{R}^{n\times n}$ will not meet it. On the other hand, the set of real ${n\times n}$ matrices having rank equal to $n-1$ can be easily seen to have dimension $n^{2}-1$, and therefore one should expect a $\mathcal{C}^{k}$ path $A(t)$ of ${n\times n}$ real matrices to lose rank (from $n$ to $n-1$) at isolated points. At such points, a singular value of $A(t)$ will become zero, in which case it generally needs to be allowed to become negative to retain its smoothness (and that of the corresponding singular vector). This is why the SVD resulting from Theorem 2.3 is called signed SVD.
@@
Also the SVD of Theorem 2.4 is a signed SVD, i.e. one must allow singular values to change sign to retain overall smoothness.
@@ 2.3 Numerical implementation of the topological test
Let $\mathcal{N}\subset\mathcal{M}$ be the embedded image of a simplex of a triangulation of $\mathcal{M}$, parametrized with respect to two parameters $(s_{1},s_{2})\in\Omega$:
@@
and such that $\mathbf{x}(\cdot)$ maps $\partial\mathcal{S}$ homeomorphically onto $\partial\mathcal{N}$. From now on, both $\mathcal{S}$ and $\mathcal{N}$ will be referred to as triangles. Since $\partial\mathcal{N}$ is a Jordan curve, we can think of it as the image of a continuous map
@@
that we can take to be periodic of least positive period 1. Recall the definitions of $J$ and $g$ in (9) and (11), respectively. With the assumptions of Theorem 2.5, our task is to numerically compute the index $\mathcal{I}_{g}(\mathcal{N})$. We do so through the following procedure:
@@ 3.2 Newton-Kantorovich Theorem and non-degeneracy of zeros
then for each $r\in\left(\frac{Y}{1-Z},r_{*}\right]$, there is a unique $\tilde{X}\in B_{r}(\bar{X})$ such that $F(\tilde{X})=0$. Moreover, the vector $\tilde{X}\in\mathbb{R}^{N}$ is a non-degenerate zero of $F$, that is $D_{X}F(\tilde{X})$ is invertible.
@@
Define the Newton-like operator $T:U\to\mathbb{R}^{N}$ by $T(X)=X-AF(X)$ and note that $D_{X}T(X)=I-AD_{X}F(X)$. Since (29) holds, pick any $r\in\left(\frac{Y}{1-Z},r_{*}\right]$. Note that $\frac{Y}{1-Z}<r$ implies that $Zr+Y<r$. The idea of the proof is to show that $T:\overline{B_{r}(\bar{X})}\to B_{r}(\bar{X})$ is a contraction.
@@ 3.3 Rigorous enclosure of $\sigma(D_{x}f(x,\lambda))$
In this short section, we demonstrate how to control the spectrum $\sigma(D_{x}f(x,\lambda))$, assuming that the matrix $D_{x}f(x,\lambda)$ is diagonalizable over the complex numbers, which is a generic condition. The idea is to introduce a pseudo-diagonalization technique and then use the Gershgorin Circle Theorem to control all eigenvalues. First, assume that using Theorem 3.3, we proved the existence of $X=(x,v,w,\lambda,h,s)\in B_{r}(\bar{X})$ such that $F(X)=0$ for some small $r>0$. Denote $\bar{X}=(\bar{x},\bar{v},\bar{w},\bar{\lambda},\bar{h},0)\in\mathbb{R}^{4n+4}$. Note that $D_{x}f(x,\lambda)$, which is a matrix whose entries can be rigorously enclosed using the control that we have over the location of $(x,\lambda)$. For instance, if the norm in $\mathbb{R}^{4n+4}$ is the sup norm $\|\cdot\|=\|\cdot\|_{\infty}$, then we know that
@@
Using interval arithmetic and the control from (30), we can construct an interval matrix $\bf M$ such that $D_{x}f(x,\lambda)\subset{\bf M}$, where the inclusion is to be understood component-wise. Then, using a standard numerical linear algebra techniques, compute a list of approximate numerical eigenvectors $\bar{\xi}_{1},\dots,\bar{\xi}_{n}\in\mathbb{C}^{n}$ and define the matrix $P$ columnwise as follows: $P=[\bar{\xi}_{1},\bar{\xi}_{2},\ldots,\bar{\xi}_{n}]\in M_{n}(\mathbb{C})$. Consider the matrix $\Lambda\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,P^{-1}D_{x}f(x,\lambda)P$, where $P^{-1}$ is the exact inverse, and hence $\sigma(D_{x}f(x,\lambda))=\sigma(\Lambda)$. Since the vectors $\bar{\xi}_{1},\dots,\bar{\xi}_{n}$ are only approximate eigenvectors, we call this process a pseudo-diagonalization. The idea now is to apply Gershgorin Circle Theorem to the matrix $\Lambda$ using the control $D_{x}f(x,\lambda)\subset{\bf M}$, that is using that
@@
While the exact computation of $D_{i}$ is in general out of reach, we can resort to interval arithmetic and use that $\Lambda\subset{\bf\Lambda}$ to construct explicit sets $\hat{D}_{i}$ such that $D_{i}\subset\hat{D}_{i}$. Hence, if we can show that the sets $\hat{D}_{i}$ are mutually disjoint, we conclude from Gershgorin Circle Theorem that each set $\hat{D}_{i}$ contains exactly one eigenvalue of $D_{x}f(x,\lambda)$.
3.4 Procedure for proving constructively existence of cusp bifurcations
@@ 4 Applications
In this section, we apply the procedure presented in this paper (as laid out explicitly in Section 3.4) to detect and prove existence of cusp bifurcations in four model problems. In all examples, we choose the sup norm $\|\cdot\|_{\infty}$ to perform the computer-assisted proofs.
@@ 4.2 Predator-prey systems with non-monotonic response function
We computed a triangulation of the portion of the manifold of equilibria for (33) contained inside the region $\{(x_{1},x_{2})\in[0,20]\times[0,20],(\lambda_{1},\lambda_{2})\in[-3,3]\times[-3,3]\}$, starting from the explicitly computed equilibrium $(x_{1},x_{2},\lambda_{1},\lambda_{2})=\left[\frac{1200}{1001},\frac{989}{1001},0,0\right]$. The computation resulted in a triangulation made of approximately 1300 nodes and 2500 simplices. See Figure 3. The detection procedure flagged 3 triangles as CUSP and 8 triangles as UNDETERMINED on step 6 of the procedure. We performed the proof on all these triangles and only one of them turned out to contain a cusp point. Further inspection revealed that the triangles that yielded a CUSP or UNDETERMINED flag were very large, a fact that happens on portions of the manifold that are locally very flat (we recall that the size of triangles is inversely proportional to local curvature). This is a warning that tolerances for the triangulation procedure should be set so to limit the largest size of triangles. In fact, tightening those tolerances resolved the issues.

### v2

@@ 1 Introduction
A cusp bifurcation is a codimension-two bifurcation that provides an organizing center at which two saddle-node bifurcations come together. This fundamental phenomenon in bifurcation theory often gives rise to a hysteresis loop, which leads to the important notion of bistability. Despite its undeniable importance in dynamical systems theory, studying rigorously the occurrence of a cusp bifurcation is hindered by two main factors. First, as a codimension-two bifurcation, detecting a cusp requires varying two parameters. Second, actually proving the existence of a cusp in a system of nonlinear differential equations is nontrivial, as it requires having a precise location of the cusp point and then evaluating carefully high order derivatives at the point. In this paper, we propose an approach which addresses these issues. To numerically detect a cusp, we perform a two-parameter continuation of equilibria (e.g. see [4, 3, 22]) along which we apply the theory of Poincaré index to search for zeros of a specific two dimensional vector field (see Section 2). Once a cusp has been numerically detected during the continuation, we apply Newton’s method to a carefully chosen map (the cusp map defined in (14)) to refine the approximation which is then validated using a computer-assisted proof based on Newton-Kantorovich theorem (see Section 3). Finally, using the explicit control of the location of the cusp, we obtain rigorously the normal form on the center manifold. To formalize the ideas, let us introduce some notation and background.
@@
The process begins with the detection step (see Section 2), which is done by computing the Poincaré index of curves (i.e. the boundaries of two-dimensional simplices used in building a triangulation of a manifold $\mathcal{M}\subset\mathcal{E}$) with respect to a carefully chosen two-dimensional vector field. The computation of the Poincaré index requires in turn to perform the continuation of the singular value decomposition of $D_{x}f(x,\lambda)$ along the boundaries of simplices. If the Poincaré index is nontrivial, then a cusp bifurcation point has been detected “inside” the simplex. This approach is different from the more standard cusp detection, which consists of the following two steps: (a) do a one-dimensional continuation (e.g. pseudo-arclength continuation, see [29]) to detect a saddle-node bifurcation (fold); and (b) free an extra parameter, perform a continuation of folds, monitor the coefficient $b$ defined in (3), and claim that a cusp bifurcation has been detected when $b$ changes sign along the branch of folds. This well-known method can be applied using standard one-dimensional continuation softwares such as AUTO [21] and MATCONT [15]. However, detecting a cusp bifurcation via one-dimensional continuations can be challenging because the chosen direction in parameter space may not go through a saddle-node bifurcation. As a result, even when starting the continuation near a cusp, this approach could fail at detecting it. In contrast, our approach relies on a two-dimensional continuation [4, 3, 22], and has the advantage of locating cusp bifurcation points on two-dimensional manifolds $\mathcal{M}\subset\mathcal{E}$. An obvious drawback is the computational overhead with respect to the approach based on one-dimensional continuations. However, a clear upside is that we construct explicitly a portion of the two-dimensional manifold $\mathcal{M}\subset\mathcal{E}$, as opposed to just one dimensional branches. In short, we get more at a larger cost.
@@ 2.2.1 Poincaré index
Let $g:(s_{1},s_{2})\in D\mapsto g(s_{1},s_{2})\in\mathbb{R}^{2}$ be continuous, with $D$ a bounded open subset of $\mathbb{R}^{2}$, and suppose $g$ has only isolated zeros in $D$. Let $\Gamma\subset D$ be a Jordan curve passing through no zeros of $g$. Let $\Delta\theta$ be the total change [^3] Let $\Gamma$ be a Jordan curve that does not contain the origin, and let it be parametrized by a periodic function $\gamma(t)$ with minimum period 1. Extend the notion of the argument of a complex number $z$, denoted by $\operatorname{Arg}(z)$, to a point on the plane in the natural way, and note that $\operatorname{Arg}$ is a multivalued map. Let $\theta$ be any continuous selection of the map $t\in[0,1]\mapsto\operatorname{Arg}(\gamma(t))$. Then, the “total change in the angle $\theta$ that $\gamma(t)$ makes with respect to any fixed direction in $\mathbb{R}^{2}$ as it traces $\Gamma$ once with positive orientation” is defined as $\theta(1)-\theta(0)$. A continuous selection exists because $\Gamma$ is compact and does not contain the origin. In fact, there are infinitely many, but the value of $\theta(1)-\theta(0)$ does not depend on the choice of the selection. in the angle $\theta$ that $g(s_{1},s_{2})$ makes with respect to some fixed direction of $\mathbb{R}^{2}$ as $(s_{1},s_{2})$ traces $\Gamma$ once with positive (i.e. counterclockwise) orientation. Then, the Poincaré index of $\Gamma$ with respect to $g$ is defined as $\Delta\theta/2\pi$, and will be denoted by $\mathcal{I}_{g}(\Gamma)$.
Clearly, the number $\mathcal{I}_{g}(\Gamma)$ is an integer. Its importance lies in the following two fundamental results, which also require $D$ to be simply connected.
@@
Let $D$ be open, bounded, and simply connected. If $\Gamma\subset D$ is a Jordan curve containing no zeros of $g$ on it or in its interior, then $\mathcal{I}_{g}(\Gamma)=0$.
@@
Let $g:(s_{1},s_{2})\in D\mapsto g(s_{1},s_{2})\in\mathbb{R}^{2}$ be continuous, where $D\subset\mathbb{R}^{2}$ is open, bounded, and simply connected. Let $\Gamma\subset D$ be a Jordan curve passing through no zeros of $g$. If $\mathcal{I}_{g}(\Gamma)\neq 0$, then $g$ has a zero inside the region bounded by $\Gamma$.
@@ 2.2.2 Singular value decomposition of matrix valued functions
It is important to realize that the assumption of Theorem 2.3 is not at all restrictive but, rather, what one should generally expect to hold for a one parameter matrix valued function. In fact, one can show (again, see [16]) that the set of real ${n\times n}$ matrices having repeated singular values has dimension $n^{2}-2$, and therefore a generic [^4] We refer to [26] for necessary background on the concept of genericity. $\mathcal{C}^{k}$ path $t\in\mathbb{R}\mapsto A(t)\in\mathbb{R}^{n\times n}$ will not meet it. On the other hand, the set of real ${n\times n}$ matrices having rank equal to $n-1$ can be easily seen to have dimension $n^{2}-1$, and therefore one should expect a $\mathcal{C}^{k}$ path $A(t)$ of ${n\times n}$ real matrices to lose rank (from $n$ to $n-1$) at isolated points. At such points, a singular value of $A(t)$ will become zero, in which case it generally needs to be allowed to become negative to retain its smoothness (and that of the corresponding singular vector). This is why the SVD resulting from Theorem 2.3 is called signed SVD.
@@
Like the SVD in Theorem 2.3, the SVD in Theorem 2.4 is also a signed SVD; that is, the singular values of $A(s_{1},s_{2})$ must be allowed to become negative to preserve their smoothness and that of the singular vectors.
@@ 2.3 Numerical implementation of the topological test
Let $\mathcal{S}$ be a simplex of a triangulation of $\mathcal{M}$, parametrized with respect to two parameters $(s_{1},s_{2})\in\Omega\subset\mathbb{R}^{2}$,
$(s_{1},s_{2})\in\Omega\mapsto\mathbf{p}(s_{1},s_{2})\in\mathcal{S},$
and let $\mathcal{N}\subset\mathcal{M}$ be the embedded [^5] One possible embedding is the orthogonal projection from $\mathcal{S}$ to $\mathcal{M}$. image of $\mathcal{S}$
@@
such that $\mathbf{x}(\cdot)$ maps $\partial\mathcal{S}$ homeomorphically onto $\partial\mathcal{N}$. From now on, both $\mathcal{S}$ and $\mathcal{N}$ will be referred to as triangles. Since $\partial\mathcal{N}$ is a Jordan curve, we can think of it as the image of a continuous map
@@
that we can take to be periodic of minimum period 1. Recall the definitions of $J$ and $g$ in (9) and (11), respectively. With the assumptions of Theorem 2.5, our task is to numerically compute the index $\mathcal{I}_{g}(\mathcal{N})$. We do so through the following procedure:
@@
  5. 5.
    In principle, one should also verify that the hypotheses of Corollary 2.2 are satisfied. However, we did not do this, as a cusp occurring on one side of $\mathcal{N}$ is clearly a non-generic event. Furthermore, such an event is automatically detected in step 7 of the detection procedure and thus addressed in point 4 of this remark.
@@ 3.2 Newton-Kantorovich Theorem and non-degeneracy of zeros
then for each $r\in\left(r_{0},r_{*}\right]$, there is a unique $\tilde{X}\in B_{r}(\bar{X})$ such that $F(\tilde{X})=0$. Moreover, the vector $\tilde{X}\in\mathbb{R}^{N}$ is a non-degenerate zero of $F$, that is $D_{X}F(\tilde{X})$ is invertible.
@@
Define the Newton-like operator $T:U\to\mathbb{R}^{N}$ by $T(X)=X-AF(X)$ and note that $D_{X}T(X)=I-AD_{X}F(X)$. Since (29) holds, pick any $r\in\left(r_{0},r_{*}\right]$. Note that $r_{0}=\frac{Y}{1-Z}<r$ implies that $Zr+Y<r$. The idea of the proof is to show that $T:\overline{B_{r}(\bar{X})}\to B_{r}(\bar{X})$ is a contraction.
@@ 3.3 Rigorous enclosure of $\sigma(D_{x}f(x,\lambda))$
In this short section, we demonstrate how to control the spectrum $\sigma(D_{x}f(x,\lambda))$, assuming that the matrix $D_{x}f(x,\lambda)$ is diagonalizable over the complex numbers, which is a generic condition. The idea is to introduce a pseudo-diagonalization technique and then use the Gershgorin Circle Theorem to control all eigenvalues. First, assume that using Theorem 3.3, we proved the existence of $X=(x,v,w,\lambda,h,s)\in B_{r}(\bar{X})$ such that $F(X)=0$ for some small $r>0$. Denote $\bar{X}=(\bar{x},\bar{v},\bar{w},\bar{\lambda},\bar{h},0)\in\mathbb{R}^{4n+4}$. Note that $D_{x}f(x,\lambda)$ is a matrix whose entries can be rigorously enclosed using the control that we have over the location of $(x,\lambda)$. For instance, if the norm in $\mathbb{R}^{4n+4}$ is the sup norm $\|\cdot\|=\|\cdot\|_{\infty}$, then we know that
@@
Using interval arithmetic and the control from (30), we can construct an interval matrix $\bf M$ such that $D_{x}f(x,\lambda)\subset{\bf M}$, where the inclusion is to be understood component-wise. Then, using a standard numerical linear algebra techniques, compute a list of approximate numerical eigenvectors $\bar{\xi}_{1},\dots,\bar{\xi}_{n}\in\mathbb{C}^{n}$ and define the matrix $P$ columnwise as follows: $P=[\bar{\xi}_{1},\bar{\xi}_{2},\ldots,\bar{\xi}_{n}]\in M_{n}(\mathbb{C})$. Consider the matrix $\Lambda\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,P^{-1}D_{x}f(x,\lambda)P$, where $P^{-1}$ is the exact inverse of the matrix $P$, and hence $\sigma(D_{x}f(x,\lambda))=\sigma(\Lambda)$. Since the vectors $\bar{\xi}_{1},\dots,\bar{\xi}_{n}$ are only approximate eigenvectors, we call this process a pseudo-diagonalization. The idea now is to apply Gershgorin Circle Theorem to the matrix $\Lambda$ using the control $D_{x}f(x,\lambda)\subset{\bf M}$, that is using that
@@
While the exact computation of $D_{i}$ is in general out of reach, we can resort to interval arithmetic and use that $\Lambda\subset{\bf\Lambda}$ to construct explicit sets $\hat{D}_{i}$ such that $D_{i}\subset\hat{D}_{i}$. Hence, if we can show that the sets $\hat{D}_{i}$ are mutually disjoint, we conclude from Gershgorin Circle Theorem (e.g. see [27]) that
$\sigma(D_{x}f(x,\lambda))\subset\bigcup_{i=1}^{n}D_{i}\subset\bigcup_{i=1}^{n}\hat{D}_{i}$
where $\hat{D}_{i}$ contains exactly one eigenvalue of $D_{x}f(x,\lambda)$.
3.4 Procedure for proving constructively the existence of cusp bifurcations
@@ 4 Applications
In this section, we apply the procedure presented in this paper (as laid out explicitly in Section 3.4) to detect and prove existence of cusp bifurcations in four model problems. In all examples, we choose the sup norm $\|\cdot\|_{\infty}$ to perform the computer-assisted proofs. Although other norms could be used in principle, the choice of the sup norm is motivated by two main factors. First, its unit ball is a hypercube aligned with the coordinate axes, making it simpler to describe and apply in various contexts. Second, when verifying the assumptions of Theorem 3.3, calculating the bound $Z$ in (28) involves computing a matrix norm. This is straightforward with the sup norm but becomes significantly more computationally intensive with other norms, such as the Euclidean norm, which requires finding the largest singular value.
@@ 4.2 Predator-prey systems with non-monotonic response function
We computed a triangulation of the portion of the manifold of equilibria for (33) contained inside the region $\{(x_{1},x_{2})\in[0,20]\times[0,20],(\lambda_{1},\lambda_{2})\in[-3,3]\times[-3,3]\}$, starting from the explicitly computed equilibrium $(x_{1},x_{2},\lambda_{1},\lambda_{2})=\left[\frac{1200}{1001},\frac{989}{1001},0,0\right]$. The computation resulted in a triangulation made of approximately 1300 nodes and 2500 simplices. See Figure 3. The detection procedure flagged 3 triangles as CUSP and 8 triangles as UNDETERMINED on step 6 of the procedure. We performed the proof on all these triangles and only one of them turned out to contain a cusp point. Further inspection revealed that the triangles that yielded a CUSP or UNDETERMINED flag were very large, a fact that happens on portions of the manifold that are locally very flat (we recall that the size of triangles is inversely proportional to local curvature). This is a warning that tolerances for the triangulation procedure should be set so to limit the size of triangles. After tightening the tolerances to reduce triangle sizes and reapplying the method, only one triangle was flagged as CUSP, and none were flagged as UNDETERMINED.